A 50-Ω resistor is connected across a battery, and a 26-mA current flows. When the 50-Ω resistor is replaced with a 23-Ω resistor, a 43-mA current flows.

A) What is the battery's voltage?

B) What is internal resistance?

I need the correct answers to 2 significant figures.
Answer #1

here by using ohms law

V = I * R

V = 26 * (R + 50)
V = 43 * (R + 23)

26 * (R + 50) = 43 * (R + 23)

17 *R = 311

R = 18.3 ohms

then by putting the values of R

V = 26 * ( 18.3 + 50)

V = 1775.8 * 10^-3 volts